OpenStudy (happykiddo):

If the remainders when three consecutive integers are each divided by 5 are 2, 3, and 4 respectively, which of the following could be the three integers? a.) 3, 4, 5 b.) 5, 6, 7 c.) 6, 7, 8 d.) 7, 8, 9 e.) 8, 9, 10 I don't understand please explain

OpenStudy (anonymous):

let the integers be x,x+1,x+2 x=5m+2 x+1=5m+3,x=5m+2 x+2=5m+4,x=5m+2 here m is an integer. when m=1 x=5*1+2=7 x+1=5*1+3=8 x+2=5*1+4=9

OpenStudy (radar):

@surjithayer Thanks for such a straight-forward solution, I was really making it complicated. Your logic is so clear. Note. It wasn't my problem, but I like to give em a try.